OpenStudy (anonymous):

the sum of a number and its positive square root is 6/25.find the number?

OpenStudy (anonymous):

x + x^1/2 = 6/25 solve for x

OpenStudy (anonymous):

we know that the value has to be lower than 1

OpenStudy (anonymous):

let the number be x. x+ sqrt(x) = 6/25 now again, let sqrt(x) be y, y^2+y = 6/25 y^2 +y - 6/25= 0 25y^2 +25y -6 = 0 factorise it and you get, (5y+6)(5y-1) = 0 y= -6/5 or y= 1/5 Since the question requires the positive square root, we will reject y= -6/5 since y=1/5 x= y^2 = 0.2^2= 0.04 The number is 0.04